BREAKING: Target Test Prep releases Brand New 2026 On Demand GMAT prep course

Redeem

Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

★★★★★5.0563 reviews
Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep
NEWSWEEK READERS’ CHOICE AWARDS 2026

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

GMATLiveTeach Starts Oct 25
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Oct 25 · Josh Braslow
Sun · 1:00 to 4:00 PM ET
Nov 8 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemand 5 seats left Start anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

Logan Thompson
EXECUTIVE ASSESSMENT

Sep 6 to Dec 6, 2026

with Logan Thompson

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemand Start anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Compare the format, schedule, and included access before enrolling. Prices and seat counts shown reflect the supplied offer details.

PROBALITY-3

Expert replies
by \'manpreet singh » Thu Dec 06, 2012 11:20 pm
Q.1) In a bag of marbles, there are 3 red, 2 white, and 5 blue. If Bob takes 2 marbles out of
the bag, what is the probability that he will have one white and one blue marble?
Join the discussion
Source: — Problem Solving |

by GMATGuruNY » Fri Dec 07, 2012 3:51 am
'manpreet singh wrote:Q.1) In a bag of marbles, there are 3 red, 2 white, and 5 blue. If Bob takes 2 marbles out of
the bag, what is the probability that he will have one white and one blue marble?
P(1st marble is white) = 2/10. (Of the 10 marbles, 2 are white.)
P(2nd marble is blue) = 5/9. (Of the 9 remaining marbles, 5 are blue.)
Since we want both events to happen, we multiply the probabilities:
2/10 * 5/9 = 1/9.
Since a good outcome will be achieved if the order of the colors is reversed, we multiply by 2:
2 * 1/9 = 2/9.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by ritind » Sat Dec 08, 2012 2:49 am
I dont understand why order is considered while taking out the probability.
Join the discussion

by GMATGuruNY » Sat Dec 08, 2012 4:07 am
ritind wrote:I dont understand why order is considered while taking out the probability.
In the problem above, there is a SINGLE EVENT: the selecting of two marbles.
We need to account for ALL OF THE WAYS that this single event could yield a white marble and a blue marble.
There are TWO WAYS that this single event could yield a favorable outcome:
Case 1: If the first marble selected is white, and the second marble selected is blue.
Case 2: If the first marble selected is blue, and the second marble selected is white.
The probability of a favorable outcome is THE SAME in each case: 1/9.
Thus, the probability that this SINGLE EVENT OF SELECTING TWO MARBLES will yield a white marble and a blue marble = 2 * 1/9 = 2/9.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by ziko » Tue Dec 11, 2012 5:38 am
One of the ways to solve this problem will be to calculate total number of combinations of having 2 marbles out of 10.
10!/2!(10-2)!=45 so there are 45 possible combinations

Now we need to find number of combinations when we have one white and one blue marble.
one white marble - 2!/1!(2-1)!=2
one blue marble - 5!/1!(5-1)!=5
So the total combinations of having 1 white and 1 blue combinations = 2*5=10

Probability of having 1 white and 1 blue = 10/45=2/9

Hope that helps!
Join the discussion